Convert 52° to radians,leaving your answer in term of pai(22/7).use your answer to find the area of sector of a circle of radius 16cm and angle 52°,giving your answer correct to two significant figures.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Convert 52° to radians,leaving your answer in term of pai(22/7)
In terms of pi, not 22/7:
52 degs = 52*pi/180 = 13pi/45 radians.
-----------------
use your answer to find the area of sector of a circle of radius 16cm and angle 52°,giving your answer correct to two significant figures.
Area = pi*r^2*(52/360)
=~ 116.17 sq cm
======================
Using radians:
Area = pi*r^2*(13pi/45)/2pi
= r^2*angle/2
= 128*(13pi/45)
= same answer.
========================
If you have the angle in radians, eg, 3 radians,
Area = r^2*3/2, slightly simpler than in degrees.

Answer by stanbon(75887)

Convert 52° to radians,leaving your answer in term of pai(22/7).use your answer to find the area of sector of a circle of radius 16cm and angle 52°,giving your answer correct to two significant figures.
----
52 degrees = 52(pi/180) = (13/45)pi radians
------------
Total area of the circle = pi*r^2 = 256pi cm^2
------------
Area of the sector = [(13pi/45)](256pi) = (74pi^2)cm^2
-----------------------------------
Cheers,
Stan H.
---------------